OpenStudy (anonymous):

hey what is tan(45+a) equal to?

OpenStudy (anonymous):

how does this (1 - tan^2a)/(1+tan^2a) cos 2a?

OpenStudy (anonymous):

(1 - tan^2a)/(1+tan^2a) = cos 2a?

OpenStudy (anonymous):

\[ \tan (a + b) = \frac{\tan a + \tan b}{1 - \tan a \tan b}\] Hmm that's the formula. One thing you can do is use this \(\large \tan a = \frac{\sin a}{\cos a}\) \[\frac{1 - \frac{\sin^2a}{\cos^2a}}{1 + \frac{\sin^2a}{\cos^2a}}\]I hope you can do the algebra from here.

OpenStudy (anonymous):

ishaan do we expand the 1 into sin^2a +cos^2a?

OpenStudy (anonymous):

yeah! :-D

OpenStudy (anonymous):

can you expand this out cause i don't really get it

OpenStudy (anonymous):

okay, i will try. \[\large \frac{\frac{cos^2a - \sin^2a}{\cos^2a}}{\frac{\cos^2a+\sin^2a}{\cos^2a}}\] \[\Large \frac{\frac{cos^2a - \sin^2a}{\cancel\cos^2a}}{\frac{1}{\cancel \cos^2a}}\] \[\cos^2a - \sin^2a = \cos 2a\]

OpenStudy (anonymous):

oh i see thank you very much

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

hey ishaan don't we get 2cos2a then and not 2/cos2a

OpenStudy (anonymous):

no i think you forgot the real equation it's \(2 \frac{1 +tan^2a}{1 - tan^2a}\)